# Bidiagonal matrix factorisations related to multiple orthogonal polynomials

### v1

@@ Abstract.
In this work, we develop two methods to obtain a factorisation as a product of bidiagonal matrices for the Hessenberg recurrence matrix of a system of multiple orthogonal polynomials. One method is based on the Gauss-Borel factorisation of the moment matrix of the system and of its Darboux transformations, and the other uses the connection of multiple orthogonal polynomials with production matrices and branched continued fractions. As a case study, we present an explicit bidiagonal factorisation for the Hessenberg recurrence matrix of the Jacobi-Piñeiro polynomials, using both methods, and of the multiple Laguerre polynomials of first kind, as a limiting case.
@@
Multiple orthogonal polynomials, Hessenberg recurrence matrix, bidiagonal factorisation, Darboux transformations, branched continued fractions, production matrices, Jacobi-Piñeiro polynomials.
@@ 1. Introduction
$\mathrm{U}=\begin{bmatrix}\alpha_{0}&1\\ &\alpha_{r+1}&1\\ &&\alpha_{2(r+1)}&1\\ &&&\ddots&\ddots\end{bmatrix}\quad\text{and}\quad\mathrm{L}_{k}=\begin{bmatrix}1\\ \alpha_{k}&1\\ &\alpha_{r+1+k}&1\\ &&\ddots&\ddots\end{bmatrix}\;\text{for}\;1\leq k\leq r,$ (1.2)
@@
whose nontrivial entries $\alpha_{n}$, $n\in\mathbb{N}$, are all different from zero.
@@
Particularly interesting are the cases where these entries are all positive, because then the Hessenberg matrix $\mathrm{H}=\mathrm{L}_{1}\cdots\mathrm{L}_{r}\,\mathrm{U}$ is oscillatory and it is the recurrence matrix of a system of multiple orthogonal polynomials with respect to positive measures on $\mathbb{R}^{+}$ (see the Favard-type theorem in [10, Thm. 6]) such that the zeros of the type II polynomials on the step-line are all simple, real, and positive, and the zeros of consecutive polynomials interlace (see [17, Thm. 4.8]). Moreover, these positive bidiagonal factorisations can be associated to Markov chains whose stochastic matrices can be factorised as products of bidiagonal stochastic matrices describing pure birth and death Markov chains (see [8, 10]).
In the literature we find examples of bidiagonal factorisations for the recurrence matrices of the Jacobi-Piñeiro multiple orthogonal polynomials (see [3, §8]) and the Hahn polynomials (see [6]), in both cases for $2$ measures, and for the recurrence matrix of the multiple orthogonal polynomials with hypergeometric moment generating functions in [17], for an arbitrary number of measures. We also mention [9], where it is obtained a bidiagonal factorisation, for $r=2$, in terms of the values of the type I and II multiple orthogonal polynomials evaluated at the origin, corresponding to Darboux and Christoffel transformations.
The main goals of this paper are to develop two methods to obtain a bidiagonal factorisation (1.1) of the Hessenberg recurrence matrix $\mathrm{H}$ of a system of multiple orthogonal polynomials, with $\alpha_{n}\neq 0$ for all $n\in\mathbb{N}$, and to establish necessary and sufficient conditions for the existence of such factorisation. In fact, this bidiagonal factorisation is not possible for every system of multiple orthogonal polynomials: an important counterexample is given when the Hessenberg matrix has only two extreme nonzero diagonals. This type of multiple orthogonal polynomials were firstly investigated in [3].
One of the methods, developed in Section 3, is based on the characterisation of multiple orthogonality via the Gauss-Borel factorisation of the corresponding moment matrix introduced in [2]. The other method, presented in Section 4, is obtained using the connection of multiple orthogonal polynomials with production matrices and branched continued fractions, introduced in [22]. In Section 5, we present an important case study of our methods: we use both algorithms to find an explicit bidiagonal factorisation of the Hessenberg recurrence matrix of the Jacobi-Piñeiro polynomials with respect to an arbitrary number of weights and we also obtain the bidiagonal factorisation of the recurrence matrix of the multiple Laguerre polynomials of first kind as a limiting case of the Jacobi-Piñeiro polynomials.
2. Background: multiple orthogonal polynomials
@@
Multiple orthogonal polynomials (see [18, § 4], [15, § 23], and [23, §3]) are a generalisation of orthogonal polynomials satisfying orthogonality conditions with respect to systems of several measures rather than just one measure. In this section, we briefly revisit the theory of multiple orthogonal polynomials, defining the orthogonality conditions via linear functionals in $\mathbb{C}[x]$, the ring of polynomials in one variable with complex coefficients. For more generality, $\mathbb{C}$ could be replaced by a generic field in all definitions and theoretical results. That generality is not relevant for this work, but it may be of interest to some readers.
A linear functional $u$ defined on $\mathbb{C}[x]$ is a linear map $u:\mathbb{C}[x]\to\mathbb{C}$. The action of a linear functional $u$ on a polynomial $f\in\mathbb{C}[x]$ is denoted by $u\left[f\right]$. The moment of order $n\in\mathbb{N}$ of a linear functional $u$ is equal to $u\left[x^{n}\right]$. By linearity, every linear functional $u$ is uniquely determined by its moments.
@@
In most relevant examples, there exists a measure $\mu$ supported on a subset of the real line or on a curve in the complex plane such that $u\left[f\right]=\int f(x)\mathrm{d}\mu(x)$ for any $f\in\mathbb{C}[x]$. However, the use of functionals provides an easier notation and allows us to not have to concern about the existence of measures.
@@
We focus on the multi-indices on the step-line: $\vec{n}=\left(n_{1},\cdots,n_{r}\right)\in\mathbb{N}^{r}$ such that $n_{1}\geq n_{2}\geq\cdots\geq n_{r}\geq n_{1}-1$. For $n\in\mathbb{N}$, there is a unique multi-index of norm $n$ on the step-line of $\mathbb{N}^{r}$. If $n=rk+j$, with $k\in\mathbb{N}$ and $10\leq j\leq r-$ 1, that multi-index is $\vec{n}=(\underbrace{k+1,\cdots,k+1}_{\text{j}},\underbrace{k,\cdots,k}_{\text{r-j}})\in\mathbb{N}^{r}$.
@@
Similarly, we consider the sequences of type I multiple orthogonal polynomials $\big{(}A_{n,1}(x),\cdots,A_{n,r}(x)\big{)}_{n\in\mathbb{N}}$ and type I linear forms $\left(q_{n}\right)_{n\in\mathbb{N}}$ on the step-line. However, for any system $\left(v_{1},\cdots,v_{r}\right)$, the conditions on the degrees of the type I multiple orthogonal polynomials for the multi-index $\vec{0}=\left(0,\cdots,0\right)\in\mathbb{N}^{r}$ imply that $A_{\vec{0},j}(x)=0$ for all $1\leq j\leq r$ and, consequently, $q_{\vec{0}}=0$. Therefore, we use the notation $A_{n,j}(x)$ and $q_{n}$, with $n\in\mathbb{N}$ and $1\leq j\leq r$, for the type I multiple orthogonal polynomials and type I linear forms for the multi-index on the step-line with norm $n+1$, so that the sequences $\big{(}A_{n,1}(x),\cdots,A_{n,r}(x)\big{)}_{n\in\mathbb{N}}$ and $\left(q_{n}\right)_{n\in\mathbb{N}}$ do not start with trivial zero elements. Using this notation:
  1. $\bullet$
@@ 2. $\bullet$
    This duality corresponds to the biorthogonality conditions in [15, Thm. 23.1.6] reduced to the step-line.
@@
with $\gamma_{n}^{\,[r]}\neq 0$ for all $n\in\mathbb{N}$ and the initial condition $P_{0}(x)=1$. The recurrence relation (2.7) satisfied by a $r$-orthogonal polynomial sequence $\left(P_{n}(x)\right)_{n\in\mathbb{N}}$ can be expressed as a vector recurrence relation
@@
$\mathrm{H}=\begin{bmatrix}\gamma^{[0]}_{0}&1\vspace*{0,1 cm}\\ \gamma^{[1]}_{0}&\gamma^{[0]}_{1}&1\vspace*{0,1 cm}\\ \vdots&\ddots&\ddots&\ddots\vspace*{0,1 cm}\\ \gamma^{[r]}_{0}&\cdots&\gamma^{[1]}_{r-1}&\gamma^{[0]}_{r}&1\\ &\ddots&&\ddots&\ddots\end{bmatrix}.$ (2.9)
Furthermore, $P_{n}(x)=\det\left(x\,\mathrm{I}_{n}-\mathrm{H}_{n}\right)$ for all $n\geq 1$.
We are interested in bidiagonal matrix factorisations of the form $\mathrm{H}=\mathrm{L}_{1}\cdots\mathrm{L}_{r}\,\mathrm{U}$ as in (1.1) for the Hessenberg matrix $\mathrm{H}$ in (2.9). Observe that this bidiagonal matrix factorisation can be used to explicitly write the coefficients of the recurrence relation (2.7) satisfied by $\left(P_{n}(x)\right)_{n\in\mathbb{N}}$ as sums of products of the nontrivial entries of the bidiagonal matrices $\mathrm{L}_{1},\cdots,\mathrm{L}_{r}$, and $\mathrm{U}$ in (1.2):
@@
with $\alpha_{m}=0$ for any $m<0$ (see [17, Thm. 4.5]). Note that there are $\binom{r+1}{k+1}$ summands in (2.10), because each summand corresponds to one way of choosing $k+1$ elements from $\{0,\cdots,r\}$.
3. Factorisations linked to the Gauss-Borel factorisation
In this section, we develop an algorithm to construct bidiagonal factorisations of the form (1.1) for the Hessenberg recurrence matrix of a system of multiple orthogonal polynomials based on the Gauss-Borel factorisation of the moment matrix of the system and of its Darboux transformations. First (in Subsections 3.1 and 3.2), we revisit results from [2] about the characterisation of mixed multiple orthogonal polynomials and their recurrence matrix via the Gauss-Borel factorisation of the moment matrix, adapting them to the setting of this paper. Then, in Subsection 3.3, we obtain the main results of this section, as we use Darboux transformations of systems of multiple orthogonal polynomials to obtain factorisations for the Hessenberg recurrence matrices as products of bidiagonal matrices, whose nontrivial entries can be expressed as coefficients of the corresponding type I and type II multiple orthogonal polynomials.
3.1. Gauss-Borel factorisation and multiple orthogonality
The moment matrix of a system of linear functionals $\left(v_{1},\cdots,v_{r}\right)$ is the infinite matrix $\mathrm{M}=\left(m_{n,\ell}\right)_{n,\ell\in\mathbb{N}}$ with entries $m_{n,rk+j}=v_{j+1}\left[x^{k+n}\right]$ for all $k,n\in\mathbb{N}$ and $10\leq j\leq r-$ 1. More explicitly,
$\mathrm{M}=\begin{bmatrix}v_{1}\left[1\right]&\cdots&v_{r}\left[1\right]&v_{1}\left[x\right]&\cdots&v_{r}\left[x\right]&v_{1}\left[x^{2}\right]&\cdots&v_{r}\left[x^{2}\right]&\cdots\vspace*{0,1 cm}\\ v_{1}\left[x\right]&\cdots&v_{r}\left[x\right]&v_{1}\left[x^{2}\right]&\cdots&v_{r}\left[x^{2}\right]&v_{1}\left[x^{3}\right]&\cdots&v_{r}\left[x^{3}\right]&\cdots\vspace*{0,1 cm}\\ v_{1}\left[x^{2}\right]&\cdots&v_{r}\left[x^{2}\right]&v_{1}\left[x^{3}\right]&\cdots&v_{r}\left[x^{3}\right]&v_{1}\left[x^{4}\right]&\cdots&v_{r}\left[x^{4}\right]&\cdots\\ \vdots&&\vdots&\vdots&&\vdots&\vdots&&\vdots&\end{bmatrix}.$ (3.1)
The Gauss-Borel factorisation of the moment matrix $\mathrm{M}$ is the LU-decomposition $\mathrm{M}=\mathrm{A}\mathrm{B}$, where $\mathrm{A}$ is a unit-lower-triangular matrix and $\mathrm{B}$ is a non-singular upper-triangular matrix, which is unique if it exists.
For $n\geq 1$, we denote by $\Delta_{n}$ the determinant of the $n$-by-$n$ leading principal submatrix of $\mathrm{M}$. All type I and type II multiple orthogonal polynomials on the step-line with respect to $\left(v_{1},\cdots,v_{r}\right)$ exist and are unique if and only if $\Delta_{n}\neq 0$ for all $n\geq 1$, because $\Delta_{n}$ is the determinant of the matrix of the system given by the multiple orthogonality conditions for the multi-index of norm $n$ on the step-line. This condition is also equivalent to the moment matrix $\mathrm{M}$ admitting a Gauss-Borel factorisation.

### v2

@@ Abstract.
We provide necessary and sufficient conditions for the Hessenberg recurrence matrix associated with a system of multiple orthogonal polynomials to admit a factorisation as a product of bidiagonal matrices. Using the Gauss-Borel factorisation of the moment matrix, we show that the nontrivial entries of those bidiagonal matrices can be expressed in terms of coefficients of type I or type II multiple orthogonal polynomials on the step-line with respect to the original system and its Christoffel transformations. Using the connection of multiple orthogonal polynomials with branched continued fractions, we show that the nontrivial entries of the bidiagonal matrices in the factorisation of the Hessenberg recurrence matrix correspond to the coefficients of a branched continued fraction associated with the given system of multiple orthogonal polynomials. As a case study, we present an explicit bidiagonal factorisation for the Hessenberg recurrence matrices of the Jacobi-Piñeiro polynomials and, as a limiting case, the multiple Laguerre polynomials of first kind.
@@
Multiple orthogonal polynomials, Hessenberg recurrence matrix, bidiagonal matrices, Gauss-Borel factorisation, branched continued fractions, Jacobi-Piñeiro polynomials.
@@ 1. Introduction and motivation
$\mathrm{L}_{i}=\begin{bmatrix}1\\ \alpha_{i}&1\\ &\alpha_{r+1+i}&1\\ &&\ddots&\ddots\end{bmatrix},\;1\leq i\leq r,\quad\text{and}\quad\mathrm{U}=\begin{bmatrix}\alpha_{0}&1\\ &\alpha_{r+1}&1\\ &&\alpha_{2(r+1)}&1\\ &&&\ddots&\ddots\end{bmatrix}$, (1.2)
@@
whose nontrivial entries $\left(\alpha_{m}\right)_{m\in\mathbb{N}}$ are all different from zero.
@@
When these entries are all positive, the Hessenberg matrix $\mathrm{H}=\mathrm{L}_{1}\cdots\mathrm{L}_{r}\,\mathrm{U}$ is oscillatory. Spectral properties of oscillatory matrices lead to a Favard-type theorem for Hessenberg matrices that admit a positive bidiagonal factorisation, in terms of a system of multiple orthogonal polynomials with respect to positive measures on $\mathbb{R}^{+}$. For these systems, the zeros of the type II polynomials on the step-line are all simple, real, and positive, and the zeros of consecutive polynomials interlace (see [10, Thm. 6] and [17, Thm. 4.8]).
The study of these positive bidiagonal factorisations dates back to [4], where high-order nonsymmetric difference operators are considered. The bidiagonal factorisation arises in connection with the vector Stieltjes continued fraction and serves as a tool for analysing both the direct and inverse spectral problems associated with the operator. It is also applied to the integration of the discrete KdV equation hierarchy.
The Hessenberg matrix arising from the recurrence relations of multiple orthogonal polynomials can be interpreted as the transition probability matrix of a Markov chain that extends beyond the classical birth-and-death framework. These positive bidiagonal factorisations correspond to Markov chains whose stochastic matrices can be decomposed into products of bidiagonal stochastic matrices, each representing pure birth-and-death Markov chains (see [8, 10]). An example of Markov chains with a physical motivation are urn models. Some urn models have been solved in terms of classical orthogonal polynomials and, more recently, a new urn model based on the bidiagonal factorisation of the Jacobi Piñeiro multiple orthogonal polynomials with respect to two measures was proposed in [15].
In the literature we find several examples of bidiagonal factorisations for the Hessenberg recurrence matrices of multiple orthogonal polynomials with respect to systems of $2$ measures. In particular, in [4, §8], we can find this bidiagonal factorisation for the Jacobi-Piñeiro polynomials, and, in [7], we can find it for the Hahn multiple orthogonal polynomials and their descendants in the multiple Askey scheme, including the Jacobi-Piñeiro polynomials, multiple Meixner (of the first and second kind), multiple Laguerre (of the first and second kind), multiple Kravchuk, and multiple Charlier. Nevertheless, the bidiagonal factorisation of the Hahn multiple orthogonal polynomials was obtained only after substantial computational effort and required an independent proof, as no general algorithm was available. Furthermore, in [9] is obtained a bidiagonal factorisation in terms of the values of the type I and II multiple orthogonal polynomials evaluated at the origin, corresponding to Darboux transformations of the Hessenberg matrix, for the Jacobi-Piñeiro and hypergeometric multiple orthogonal polynomials, again with respect to $2$ measures.
Bidiagonal factorisations of the Hessenberg recurrence matrices associated with multiple orthogonal polynomials with respect to systems of any number of measures whose moment generating functions are hypergeometric were obtained in [17], using a connection of those polynomials with branched continued fractions. That paper provides an explicit example of a new family of multiple orthogonal polynomials, for which the location, simplicity, and interlacing of the zeros were proved using the bidiagonal factorisation of the corresponding Hessenberg recurrence matrix and there is no other known method to prove these properties without imposing significantly stronger conditions on the parameters involved.
The main goals of this paper are to develop methods to obtain bidiagonal factorisations of the form (1.1) for the Hessenberg recurrence matrices of systems of multiple orthogonal polynomials and to establish necessary and sufficient conditions for the existence of those factorisations. In Section 3, we give a sufficient condition for the existence of the bidiagonal factorisation. In this approach, the nontrivial entries $\left(\alpha_{n}\right)_{n\in\mathbb{N}}$ in the bidiagonal matrices are described in terms of the coefficients of the type I or type II polynomials along the step-line, for both the original system and its Christoffel transformations. In Section 4, we obtain the same sequence $\left(\alpha_{n}\right)_{n\in\mathbb{N}}$ using the connection between multiple orthogonal polynomials, production matrices and branched continued fractions as introduced in [24]. This framework allows us to establish necessary and sufficient conditions for the existence of the bidiagonal factorisation. Finally, in Section 5, we present a case study: we construct the explicit bidiagonal factorisation of the Hessenberg recurrence matrix of the Jacobi-Piñeiro polynomials with an arbitrary number of weights. As a limiting case, we also derive the corresponding factorisation for the multiple Laguerre polynomials of first kind.
2. Background
2.1. Multiple orthogonal polynomials
@@
Multiple orthogonal polynomials (see [20, § 4], [16, § 23], and [26, §3]) are a generalisation of orthogonal polynomials satisfying orthogonality conditions with respect to systems of several measures rather than just one measure. In this section, we briefly revisit the theory of multiple orthogonal polynomials, defining the orthogonality conditions via linear functionals in $\mathbb{C}[x]$, the ring of polynomials in one variable with complex coefficients. For more generality, $\mathbb{C}$ could be replaced by a generic field in all definitions and theoretical results. That generality is not relevant for this work, but it may be of interest to some readers.
@@
A linear functional $u$ defined on $\mathbb{C}[x]$ is a linear map $u:\mathbb{C}[x]\to\mathbb{C}$. The action of a linear functional $u$ on a polynomial $f\in\mathbb{C}[x]$ is denoted by $u\left[f\right]$. The moment of order $n\in\mathbb{N}$ of a linear functional $u$ is equal to $u\left[x^{n}\right]$. By linearity, every linear functional $u$ is uniquely determined by its moments. In most relevant examples, there exists a measure $\mu$ supported on a subset of the real line or on a curve in the complex plane such that $u\left[f\right]=\int f(x)\mathrm{d}\mu(x)$ for any $f\in\mathbb{C}[x]$ and we say that the functional $u$ is induced by the measure $\mu$. However, the use of functionals allows us to not have to concern about the existence of measures in most theoretical results (and it also provides a shorter notation).
@@
In most relevant examples, the functionals $v_{1},\cdots,v_{r}$ are induced by measures $\mu_{1},\cdots,\mu_{r}$ which are all absolutely continuous with respect to a common measure $\mu$, i.e., there exist weight functions $w_{1}(x),\cdots,w_{r}(x)$ defined in the support of $\mu$ such that $v_{j}\left[f\right]=\int f(x)w_{j}(x)\mathrm{d}\mu(x)$ for any $f\in\mathbb{C}[x]$ and any $1\leq j\leq r$. Then, the type I function $Q_{\vec{n}}(x)$ is defined by $Q_{\vec{n}}(x)=\sum_{j=1}^{r}A_{\vec{n},j}(x)\,w_{j}(x)$ and the type I linear form $q_{\vec{n}}$ defined in (2.2) has the integral representation $q_{\vec{n}}\left[f\right]=\int f(x)Q_{\vec{n}}(x)\mathrm{d}\mu(x)$ for any $f\in\mathbb{C}[x]$.
@@
We focus on the multi-indices on the step-line: $\vec{n}=\left(n_{1},\cdots,n_{r}\right)\in\mathbb{N}^{r}$ such that $n_{1}\geq n_{2}\geq\cdots\geq n_{r}\geq n_{1}-1$. For $n\in\mathbb{N}$, there is a unique multi-index of norm $n$ on the step-line of $\mathbb{N}^{r}$. If $n=rk+j$, with $k\in\mathbb{N}$ and $0\leq j\leq r-1$, that multi-index is formed by $j$ entries equal to $k+1$ followed by rjr-j entries equal to $k$.
@@
Similarly, we consider the sequences of type I multiple orthogonal polynomials $\big(A_{n,1}(x),\cdots,A_{n,r}(x)\big)_{n\in\mathbb{N}}$ and type I linear forms $\left(q_{n}\right)_{n\in\mathbb{N}}$ on the step-line. However, for any system $\left(v_{1},\cdots,v_{r}\right)$, the conditions on the degrees of the type I multiple orthogonal polynomials for the multi-index $\vec{0}=\left(0,\cdots,0\right)\in\mathbb{N}^{r}$ imply that $A_{\vec{0},j}(x)=0$ for all $1\leq j\leq r$ and, consequently, $q_{\vec{0}}=0$.
Therefore, we use the notation $A_{n,j}(x)$ and $q_{n}$, with $n\in\mathbb{N}$ and $1\leq j\leq r$, for the type I multiple orthogonal polynomials and type I linear forms for the multi-index on the step-line with norm $n+1$, so that the sequences $\big(A_{n,1}(x),\cdots,A_{n,r}(x)\big)_{n\in\mathbb{N}}$ and $\left(q_{n}\right)_{n\in\mathbb{N}}$ do not start with trivial zero elements. Using this notation:
@@
    This duality corresponds to the biorthogonality conditions in [16, Thm. 23.1.6] reduced to the step-line.
@@
with $\gamma_{n}^{\,[r]}\neq 0$ for all $n\in\mathbb{N}$ and the trivial initial condition $P_{0}(x)=1$. The recurrence relation (2.7) satisfied by a $r$-orthogonal polynomial sequence $\left(P_{n}(x)\right)_{n\in\mathbb{N}}$ can be expressed as a vector recurrence relation
@@
$\mathrm{H}=\begin{bmatrix}\gamma^{[0]}_{0}&1\vskip 2.84544pt\\ \gamma^{[1]}_{0}&\gamma^{[0]}_{1}&1\vskip 2.84544pt\\ \vdots&\ddots&\ddots&\ddots\vskip 2.84544pt\\ \gamma^{[r]}_{0}&\cdots&\gamma^{[1]}_{r-1}&\gamma^{[0]}_{r}&1\\ &\ddots&&\ddots&\ddots\end{bmatrix}$. (2.9)
Furthermore, for any $n\geq 1$, $P_{n}(x)$ is the characteristic polynomial of the truncated $n$-by-$n$ Hessenberg matrix $\mathrm{H}_{n}$ formed by the first $n$ rows and columns of $\mathrm{H}$. As a result, the study of the spectral properties of the matrices $\mathrm{H}_{n}$ gives information about the zeros of the $r$-orthogonal polynomials $P_{n}(x)$.
We are interested in bidiagonal matrix factorisations of the form $\mathrm{H}=\mathrm{L}_{1}\cdots\mathrm{L}_{r}\,\mathrm{U}$ as in (1.1) for the Hessenberg matrix $\mathrm{H}$ in (2.9). Observe that these bidiagonal matrix factorisations can be used to explicitly write the coefficients of the recurrence relation (2.7) satisfied by $\left(P_{n}(x)\right)_{n\in\mathbb{N}}$ as sums of products of the nontrivial entries of the bidiagonal matrices $\mathrm{L}_{1},\cdots,\mathrm{L}_{r}$, and $\mathrm{U}$ in (1.2):
@@
with $\alpha_{m}=0$ for any $m<0$ (see [17, Thm. 4.5]).
2.2. Gauss-Borel factorisation of the moment matrix and multiple orthogonality
The moment matrix of a system of linear functionals $\left(v_{1},\cdots,v_{r}\right)$ is the infinite matrix $\mathrm{M}=\left(m_{n,\ell}\right)_{n,\ell\in\mathbb{N}}$ with entries $m_{n,rk+j}=v_{j+1}\left[x^{k+n}\right]$ for all $k,n\in\mathbb{N}$ and $0\leq j\leq r-1$. More explicitly,
$\mathrm{M}=\begin{bmatrix}v_{1}\left[1\right]&\cdots&v_{r}\left[1\right]&v_{1}\left[x\right]&\cdots&v_{r}\left[x\right]&v_{1}\left[x^{2}\right]&\cdots&v_{r}\left[x^{2}\right]&\cdots\vskip 2.84544pt\\ v_{1}\left[x\right]&\cdots&v_{r}\left[x\right]&v_{1}\left[x^{2}\right]&\cdots&v_{r}\left[x^{2}\right]&v_{1}\left[x^{3}\right]&\cdots&v_{r}\left[x^{3}\right]&\cdots\vskip 2.84544pt\\ v_{1}\left[x^{2}\right]&\cdots&v_{r}\left[x^{2}\right]&v_{1}\left[x^{3}\right]&\cdots&v_{r}\left[x^{3}\right]&v_{1}\left[x^{4}\right]&\cdots&v_{r}\left[x^{4}\right]&\cdots\\ \vdots&&\vdots&\vdots&&\vdots&\vdots&&\vdots&\end{bmatrix}$. (2.11)
The Gauss-Borel factorisation of the moment matrix $\mathrm{M}$ is the LU-decomposition $\mathrm{M}=\mathrm{C}\mathrm{D}$, where $\mathrm{C}$ is a unit-lower-triangular matrix and $\mathrm{D}$ is a non-singular upper-triangular matrix, which is unique if it exists. The Gauss-Borel factorisation exists if and only if $\Delta_{n}\neq 0$ for all $n\geq 1$, where $\Delta_{n}$ is the determinant of the $n$-by-$n$ leading principal submatrix of $\mathrm{M}$. This condition is equivalent to the existence and uniqueness of the type I and type II multiple orthogonal polynomials on the step-line with respect to $\left(v_{1},\cdots,v_{r}\right)$.